# AutoVP: An Automated Visual Prompting Framework and Benchmark

### v1

@@ B.10 Exploring Additional Tuning Axes
In AutoVP, we set the learning rate (LR) for CLIP to be 40 and employed the SGD optimizer with a momentum of 0.9 and a weight decay (WD) of 0. In this section, we have undertaken supplementary tuning options within the AutoVP framework. Specifically, we introduce additional choices for the LR, selecting from 35, 40, and 45 (for CLIP), as well as WD, choosing from 0, $10^{-5}$, and $10^{-10}$. The outcomes presented in Table 8 showcase a 0.6% enhancement in accuracy with the additional tuning options. It is important to note that the execution workload dramatically increases, as it will involve exploring 9 times more additional combinations in the tuning process.
@@
Table 8: Hyper-Parameter Tuning for Learning Rate (LR) and Weight Decay (WD). The table displays the optimal tuning configurations and corresponding testing accuracy for both scenarios with and without additional LR and WD tuning options for AutoVP on Flowers102. The pre-trained model used is CLIP. The highest accuracy is marked in bold.
@@
B.11 Improved ILM-VP with Tuning Configuration
@@
The tuning procedure is applied to several datasets, and the results are shown in Table 9. Although hyper-parameter tuning in ILM-VP leads to an improvement in accuracy, it remains unable to surpass the performance exhibited by AutoVP.
Table 9: ILM-VP with Hyper-Parameter Tuning. The chosen tuning configurations and the corresponding testing accuracy (%) in ILM-VP. The highest accuracy is marked in bold.
@@
B.12 Computing Resources
In this section, we provide the execution time (measured on NVIDIA GeForce RTX 3090) and the comparison of trainable parameters of AutoVP. For ease of comparison, we use the Flower102 dataset for illustration. In Table 10, we provide the end-to-end execution time (hyper-parameter tuning + 200-epoch training) of AutoVP. When we measure only the 200-epoch training, AutoVP demonstrates its competitiveness in terms of similar or even lower training time (see Table 11), compared to the state-of-the-art VP baselines (ILM-VP and CLIP-VP The comparison of trainable parameters can also be found in Table 12.
Table 10: Flowers102 End-to-End Execution Time. For the Flowers102 dataset, the pre-trained model selected by AutoVP is Swin-T, the output mapping is FullyMap, and the prompt size is 16.
@@ B.12 Computing Resources
Time (minutes) 146 31 177
Table 11: Execution Time Comparison. The 200-epoch training time (in minutes) on the Flowers102 dataset varies depending on the chosen pre-trained model: Swin-T or CLIP. In both cases, AutoVP utilizes the FullyMap as the output mapping method with the prompt size 16.
@@
Table 12: Trainable Parameter Size. The average trainable parameter sizes (million) are calculated across the 12 datasets for different pre-trained models, mapping methods, and baselines.
AutoVP
SemanticMap
AutoVP
FreqMap
AutoVP
IterMap
AutoVP
FullyMap
Linear Probing Full Finetune

### v2

@@ E.3 Exploring Additional Tuning Axes
In AutoVP, we set the learning rate (LR) for CLIP to be 40 and employed the SGD optimizer with a momentum of 0.9 and a weight decay (WD) of 0. In this section, we have undertaken supplementary tuning options within the AutoVP framework. Specifically, we introduce additional choices for the LR, selecting from 35, 40, and 45 (for CLIP), as well as WD, choosing from 0, $10^{-5}$, and $10^{-10}$. The outcomes presented in Table E.3 showcase a 0.6% enhancement in accuracy with the additional tuning options. It is important to note that the execution workload dramatically increases, as it will involve exploring 9 times more additional combinations in the tuning process.
@@
Table E.3: Hyper-Parameter Tuning for Learning Rate (LR) and Weight Decay (WD). The table displays the optimal tuning configurations and corresponding testing accuracy for both scenarios with and without additional LR and WD tuning options for AutoVP on Flowers102. The pre-trained model used is CLIP. The highest accuracy is marked in bold.
@@
E.4 Improved ILM-VP with Tuning Configuration
@@
The tuning procedure is applied to several datasets, and the results are shown in Table E.4. Although hyper-parameter tuning in ILM-VP leads to an improvement in accuracy, it remains unable to surpass the performance exhibited by AutoVP.
Table E.4: ILM-VP with Hyper-Parameter Tuning. The chosen tuning configurations and the corresponding testing accuracy (%) in ILM-VP. The highest accuracy is marked in bold.
@@
E.5 Comparison of AutoVP and BlackVIP
Table E.5: AutoVP vs. BlackVIP. The testing accuracy (%) of AutoVP and BlackVIP.
Pets Cars Flowers Food SUN DTD SVHN EuroSAT RESISC CLEVR UCF Avg.
AutoVP 88.2 61.8 90.4 82.3 65.4 62.5 92.9 96.8 88.5 82.8 73.1 80.4
BlackVIP 89.7 65.6 70.6 86.6 64.7 45.2 44.3 73.1 64.5 36.8 69.1 64.6
Some visual prompting research has delved into a black-box setting (Tsai et al., 2020; Oh et al., 2023), where the internal architecture of the pre-trained model remains unattainable during the training process. For instance, BlackVIP (Oh et al., 2023) employs an input-dependent prompt designer to generate visual prompts. These prompts are then fed into a black-box model, and gradient approximation strategies are utilized to update the prompt designer. In Table E.5, we present a comparison between AutoVP and BlackVIP. Since BlackVIP uses CLIP as the pre-trained backbone, we report our results using the same pre-trained model. AutoVP demonstrates a 16% performance increase on average compared to BlackVIP. This notable difference might be attributed to the variance in update strategies: BlackVIP utilizes SPSA-GC for black-box models, while AutoVP relies on classic gradient descent. However, due to the differing objectives of these two studies, direct comparisons may introduce certain unfairness.
Appendix F Performance and Resource Utilization
F.1 Comparison of AutoVP, Linear Probing, and Full Fine-Tuning
When comparing the performance of AutoVP to that of linear probing (LP) and full fine-tuning (FF), it’s crucial to acknowledge the significant discrepancy in terms of trainable parameter size. As FF involves a parameter size that is roughly 100 to 1000 times larger (as shown in Table F.4). Nevertheless, it’s worth noting that the differences in performance between VP and FF are relatively minor in certain datasets. For instance, in the case of EuroSAT, both AutoVP and FF achieve a 96% accuracy (as demonstrated in Table F.1). This observation suggests that VP retains its advantages even when faced with such substantial variations in parameter size.
Table F.1: Comparison of AutoVP, Linear Probing (LP), and Full Fine-Tuning (FF). With the EuroSAT dataset and using CLIP as the pre-trained model, the table displays the testing accuracy, execution time, and trainable parameter size associated with each respective method.
Experimental Info. AutoVP LP FF
Accuracy (%) 96.84 94.62 96.78
Execution Time (second) 2448 2370 3081
Trainable Parameter Size (Million) 0.15 0.005 151.28
F.2 Computing Resources
In this section, we provide the execution time (measured on NVIDIA GeForce RTX 3090) and the comparison of trainable parameters of AutoVP. For ease of comparison, we use the Flower102 dataset for illustration. In Table F.2, we provide the end-to-end execution time (hyper-parameter tuning + 200-epoch training) of AutoVP. When we measure only the 200-epoch training, AutoVP demonstrates its competitiveness in terms of similar or even lower training time (see Table F.3), compared to the state-of-the-art VP baselines (ILM-VP and CLIP-VP The comparison of trainable parameters can also be found in Table F.4.
Table F.2: Flowers102 End-to-End Execution Time. For the Flowers102 dataset, the pre-trained model selected by AutoVP is Swin-T, the output mapping is FullyMap, and the prompt size is 16.
@@ F.2 Computing Resources
Time 146 mins 31 mins 177 mins
Table F.3: Execution Time Comparison. The 200-epoch training time (in minutes) on the Flowers102 dataset varies depending on the chosen pre-trained model: Swin-T or CLIP. In both cases, AutoVP utilizes the FullyMap as the output mapping method with the prompt size 16.
@@
Table F.4: Trainable Parameter Size. The average trainable parameter sizes (million) are calculated across the 12 datasets for different pre-trained models, mapping methods, and baselines.
AutoVP Linear Probing Full Finetune
SemanticMap FreqMap IterMap FullyMap